A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. n equals 13n=13​, p equals 0.25p=0.25​, x less than or equals 4x≤4 The probability of x less than or equals 4x≤4 successes is
Answer #1

t n=13, p= 0.25 P(x=x) pr (1-P) 2 TP(XS4) = P(X=0) + P(x-1) + P(X=2) + P(X23) -t plX24) P[X20) - (2) 0.25 0.755 = 0.0238 P(